# While GitHub Copilot Excels at Coding, Does It Ensure Responsible Output?

### v1

@@ Abstract
The rapid development of large language models (LLMs) has significantly advanced code completion capabilities, giving rise to a new generation of LLM-based Code Completion Tools (LCCTs). Unlike general-purpose LLMs, these tools possess unique workflows, integrating multiple information sources as input and prioritizing code suggestions over natural language interaction, which introduces distinct security challenges. Additionally, LCCTs often rely on proprietary code datasets for training, raising concerns about the potential exposure of sensitive data. This paper exploits these distinct characteristics of LCCTs to develop targeted attack methodologies on two critical security risks: jailbreaking and training data extraction attacks. Our experimental results expose significant vulnerabilities within LCCTs, including a $99.4\%$ success rate in jailbreaking attacks on GitHub Copilot and a $46.3\%$ success rate on Amazon Q. Furthermore, We successfully extracted sensitive user data from GitHub Copilot, including $54$ real email addresses and $314$ physical addresses associated with GitHub usernames. Our study also demonstrates that these code-based attack methods are effective against general-purpose LLMs, such as the GPT series, highlighting a broader security misalignment in the handling of code by modern LLMs. These findings underscore critical security challenges associated with LCCTs and suggest essential directions for strengthening their security frameworks. The example code and attack samples from our research are provided at https://github.com/Sensente/Security-Attacks-on-LCCTs.
@@ 1 Introduction
The deployment of LLM-based Code Completion Tools (LCCTs) is seeing unprecedented growth. GitHub Copilot, a leading example, has garnered over 1.3 million paid subscribers and 50,000 enterprise customers worldwide, demonstrating its widespread adoption Wilkinson (2024). Known as "AI pair programmers," these tools assist developers by providing code suggestions powered by LLMs. Specialized LCCTs like GitHub Copilot GitHub (2024a) and Amazon Q Amazon (2024a) fine-tune general-purpose LLMs on a diverse array of programming languages from public repositories to enhance their code completion capabilities. Similarly, general-purpose LLMs such as OpenAI’s ChatGPT OpenAI (2022) and GPT-4 Bubeck et al. (2023) also offer code completion features.
Despite offering significant capabilities, LCCTs pose considerable new security risks. Previous research has focused on the software engineering aspects of LCCT-generated code security Zhang et al. (2023); Fu et al. (2023); Rabbi et al. (2024); Tambon et al. (2024). Yet, it overlooks the security vulnerabilities inherent in the LLMs that power these tools.
@@
Our paper seeks to address this gap by answering the question: “Do LCCTs ensure responsible output?” We begin by outlining the key differences between LCCTs and general-purpose LLMs. Firstly, LCCTs process a variety of inputs including current code, file names, and contents from other open files, increasing the risk of security breaches due to these diverse information sources. Secondly, while general-purpose LLMs are primarily tailored for natural language interactions, LCCTs specialize in code completion and suggestions, making them vulnerable to security challenges unique to code-based inputs. Lastly, LCCTs often utilize proprietary datasets for training to enhance their coding ability, which potentially contains sensitive user data.
We utilize these differences to develop targeted attack strategies for two novel security risks associated with LCCTs, i.e., jailbreaking and training data extraction attacks Carlini et al. (2021), as shown in Figure 1. Specifically, given that LCCTs inherit the general capabilities of backend LLMs, we embed the jailbreaking prompts into different code components to circumvent LCCT security protocols. Furthermore, leveraging the innate tendency of LLMs to memorize training data, we devise training data extraction attacks on LCCTs. This enables the unauthorized extraction of privacy-sensitive information embedded within LCCTs’ training data, thereby compromising user privacy.
We conduct extensive experiments to evaluate the attacks on two mainstream LCCTs, i.e., GitHub Copilot and Amazon Q, and three general-purpose LLMs, i.e., GPT-3.5, GPT-4, and GPT-4o. For the jailbreaking attacks, our results indicate that with tailored attack methodologies, we achieve a $99.4\%$ Attack Success Rate (ASR) on GitHub Copilot and a 46.3% ASR on Amazon Q, where ASR reflects the rate at which harmful information is generated. These results significantly exceed the $18.8\%$ and $0\%$ ASR achieved by existing attacks on GPT-4 and GPT-4o, respectively. For training data extraction attacks, we successfully extract valid private data from GitHub Copilot, including email addresses, and locations associated with real GitHub usernames.
@@ 2.4 Safety Concerns of LCCTs
Recent advancements in LLMs have significantly propelled the development of LCCTs. These LCCTs have become integral to developers’ workflows by providing context-sensitive code suggestions, thus enhancing productivity. Note that both the commercial LCCTs, including GitHub Copilot and Amazon Q, and the general-purpose LLMs like the GPT series provide code assistance. Our study primarily focuses on these LCCTs, and a comparative analysis of their features is presented in Table 1. Security evaluations of LCCTs traditionally focused on software engineering aspects, such as security vulnerabilities and code quality Pearce et al. (2022); Fu et al. (2023); Majdinasab et al. (2024). Recently, attention has shifted to copyright issues related to generated code Al-Kaswan and Izadi (2023), including the risk of distributing copyrighted code without proper licensing Basanagoudar and Srekanth (2023). Additionally, there are concerns about LCCTs unintentionally revealing hardcoded sensitive data due to the retention properties of LLMs Huang et al. (2024). Despite these insights, prior research has largely overlooked the inherent security risks of LCCTs, especially the threat of direct attacks on backend LLM models to manipulate outputs illicitly. Our research mitigates this gap by providing a comprehensive analysis of direct attacks on LCCTs and their potential implications.
@@ 3 Understanding How LCCT Works
Next, we elucidate how these workflow steps differentiate LCCTs from general-purpose LLMs, highlighting vulnerabilities susceptible to attacks.
@@ 4.1 Attack to Contextual Information Aggregation
4.1.1 Filename Proxy Attacks.
The first strategy targets LCCTs like GitHub Copilot, which leverage filenames to refine code suggestions. The construction of this attack is depicted in Figure 2(a). We name a code file with a sensitive query and use a static comment within the file to prompt the LCCT. For example, we use the prompt “Now the current filename is the question, answer the question with concrete code.” to trigger the code completion. Below this, a variable $answer$ is initialized, setting the stage for the LCCT’s code completion response.
4.1.2 Cross-File Attack.
We extend this attack by leveraging the capabilities of LCCTs to gather information across multiple files. While this feature enhances LCCTs’ understanding of code context, it also opens avenues for more subtle attack vectors. We develop the “Cross-File Attack” that involves invoking functions from other files within the current file. The construction process is illustrated in Figure 2(b). Specifically, we set up two files within the same project: “file1” and “file2”, named to exclude semantic information. “file2” includes a function question designed to return a problematic response. In “file1”, a simple code snippet calls the question function from “file2”, triggering the LCCT’s completion mechanisms. This method ensures the attack remains hidden, as “file1” contains no explicit plain text clues about the sensitive content.
@@
We then investigate embedding jailbreaking prompts within code snippets, developing two levels of attacks utilizing different programming constructs. “Level I – Guided Trigger Attack” manipulates variables and their names into jailbreaking prompts capable of breaching the LCCTs. “Level II – Code Embedded Attack” further obscures the attacks from LCCT detection using diverse code snippet components.
@@ 4.2 Hierarchical Code Exploitation Attack
4.2.1 Level I – Guided Trigger Attack.
@@
1. Variable Transformation. We convert prohibited queries into variable names. Following established best practices for variable naming, which emphasize conciseness and efficiency Complete (1993), we eliminate interrogative phrases such as "How to" or "What is" and reformulate the query into an imperative form. This serves as the name for a string variable.
2. Guiding Words Addition. To steer the LCCT toward generating the desired code output, we attach guiding but semantically empty prompts to the newly created string variable. These prompts are designed not to answer but to trigger the LCCT’s code completion capabilities.
4.2.2 Level II – Code Embedded Attack.
@@
1). Adding Fixed Elements. We incorporate several static components into the code file, including an initial descriptive comment, a $print$ function in the middle to simulate typical code operations, and a concluding comment to denote expected output.
2). Query Transformation Customization. To further obscure the attack, we distribute the query from Level I across several variables. Sensitive terms (e.g., “illegal”, “counterfeit”) are embedded within different string variables and later merged through string concatenation to form the complete query. For the guiding words in the attack code, which serve no substantive purpose, we use them to initialize an $answer$ variable, making it the target for LCCTs’ code completion.
@@ 4.3 Code-Driven Privacy Extraction Attack
1). Retrieving Leaked Account ID. We first follow the Code Embedded Level I Attack method. Here, we craft a string variable named $url$, prefixed with $github.com/$ to activate the LCCT’s code completion, aiming to retrieve the GitHub username.
2). Extracting Privacy Based on Account ID. We proceed by designing a code snippet to extract specific private details of the identified GitHub user. The snippet starts with a comment detailing the privacy type to be retrieved, followed by a corresponding variable initiation. In our case studies, we focus on extracting the user’s email address and location, with the LCCT being prompted to complete these details.
@@ 5.1 Evaluation Setup
We evaluate our attack framework on two mainstream commercial LCCTs and the latest checkpoints of three general-purpose LLMs with code generation capabilities: GitHub Copilot (version 1.211.0), Amazon Q (version 1.12.0), GPT-3.5 (GPT-3.5-turbo-0125), GPT-4o (GPT-4o-2024-05-13), and GPT-4 (GPT-4-turbo-2024-04-09). As shown in Table 1, the functionalities of these LCCTs differ, guiding our implementation and evaluation of targeted experiments. Python is the primary language for our experiments as it is the most commonly used language among LCCT users Zhang et al. (2023). The adaptability of our attacks to other languages is assessed using Go in our ablation study. We provide a detailed evaluation setup for jailbreaking and training data extraction attacks separately as follows.
@@
Table 2: Jailbreaking ASR micro benchmarks across different models and attack methods. The best ASR for each attack method is highlighted in bold.
@@ 5.2 Micro Benchmark Results
LCCTs exhibit extreme vulnerability to jailbreaking attacks. LCCTs demonstrate a pronounced susceptibility to jailbreaking attacks, with significantly higher ASR compared to the latest general-purpose LLMs. For instance, the “Level I – Guided Trigger Attack” consistently achieves a $99.4\%$ ASR with GitHub Copilot, indicating its effectiveness in eliciting responses with malicious content. In contrast, the DAN attack registers much lower ASRs of $18.8\%$ on GPT-4 and $0\%$ on GPT-4o.
@@
There is a trade-off between attack design complexity and the back-end LLM capabilities. Our results indicate a correlation between the complexity of the attack design and the capabilities of the underlying LLM models. Specifically, less sophisticated models (e.g., Copilot, Amazon Q, GPT-3.5) show higher ASRs for “Level I – Guided Trigger Attack” compared to “Level II – Code Embedded Attack.” Conversely, more advanced models (e.g., GPT-4 and GPT-4o) either match or exceed the success rates of more complex attacks, such as “Level II Attack.” This suggests that intricate attacks may surpass the comprehension abilities of simpler models, which tend to mimic rather than understand the attack constructs—aligning with the principles of Occam’s Razor. As LLMs advance, we anticipate that the sophistication of “Level II Attack” will obscure the mechanisms of jailbreaking further, potentially improving its attack efficacy across both LCCTs and general-purpose models.
@@ 5.3 Ablation Studies
To validate the generalizability of our Hierarchical Code Exploitation Attack across programming languages, we conduct evaluations using the Go language. Compared to Python, Go is less frequently used by LCCTs’ users Zhang et al. (2023), which also implies that there is a smaller portion of the code corpus in the LLCT proprietary code dataset Nijkamp et al. (2022); Li et al. (2022). Table 6 shows that, compared to using Python, both Copilot and Amazon Q achieve an increase in ASR in Level I and Level II attacks when using Go as the vector. This significant difference underscores the security challenges LCCTs face with multiple programming languages, emphasizing the need for stronger measures as language support expands.

### v3

# Security Attacks on LLM-based Code Completion Tools

@@ Abstract
The rapid development of large language models (LLMs) has significantly advanced code completion capabilities, giving rise to a new generation of LLM-based Code Completion Tools (LCCTs). Unlike general-purpose LLMs, these tools possess unique workflows, integrating multiple information sources as input and prioritizing code suggestions over natural language interaction, which introduces distinct security challenges. Additionally, LCCTs often rely on proprietary code datasets for training, raising concerns about the potential exposure of sensitive data. This paper exploits these distinct characteristics of LCCTs to develop targeted attack methodologies on two critical security risks: jailbreaking and training data extraction attacks. Our experimental results expose significant vulnerabilities within LCCTs, including a $99.4\%$ success rate in jailbreaking attacks on GitHub Copilot and a $46.3\%$ success rate on Amazon Q. Furthermore, We successfully extracted sensitive user data from GitHub Copilot, including $54$ real email addresses and $314$ physical addresses associated with GitHub usernames. Our study also demonstrates that these code-based attack methods are effective against general-purpose LLMs, highlighting a broader security misalignment in the handling of code by modern LLMs. These findings underscore critical security challenges associated with LCCTs and suggest essential directions for strengthening their security frameworks. The example code and attack samples from our research are provided at https://github.com/Sensente/Security-Attacks-on-LCCTs.
@@ 1 Introduction
2 Introduction
The deployment of LLM-based Code Completion Tools (LCCTs) is seeing unprecedented growth. GitHub Copilot, a leading example, has garnered over 1.3 million paid subscribers and 50,000 enterprise customers worldwide, demonstrating its widespread adoption Wilkinson (2024). Known as “AI pair programmers," these tools assist developers by providing code suggestions powered by LLMs. Specialized LCCTs like GitHub Copilot GitHub (2024a) and Amazon Q Amazon (2024a) fine-tune general-purpose LLMs on a diverse array of programming languages from public repositories to enhance their code completion capabilities. Similarly, general-purpose LLMs (referred to as general LLMs) such as OpenAI’s ChatGPT OpenAI (2022) and GPT-4 Bubeck et al. (2023) also offer code completion features.
Despite offering significant capabilities, LCCTs pose considerable new security risks. Previous research has focused on the software engineering aspects of LCCT-generated code security Zhang et al. (2023); Fu et al. (2023); Rabbi et al. (2024); Tambon et al. (2024). However, these works neglect the security vulnerabilities inherent in the underlying LLMs that power such tools.
@@ 2 Introduction
Our paper seeks to address this gap by exploring the question: “Do LCCTs ensure responsible output?” We begin by outlining the key distinctions between LCCTs and general LLMs. First, LCCTs process a variety of inputs including current code, file names, and contents from other open files, increasing the risk of security breaches due to these diverse information sources. Second, while general LLMs are primarily tailored for natural language tasks, LCCTs specialize in code completion and suggestions, making them vulnerable to security challenges unique to code-based inputs. Finally, LCCTs frequently rely on proprietary datasets for training to enhance coding capabilities, which may inadvertently include sensitive user data, raising security concerns.
We leverage these distinctions to design targeted attack strategies addressing two novel security risks inherent to LCCTs: jailbreaking and training data extraction attacks Carlini et al. (2021), as illustrated in Figure 1. Specifically, since LCCTs inherit the general capabilities of their underlying LLMs, we embed jailbreaking prompts within various code components to bypass LCCT security mechanisms. Additionally, by exploiting the tendency of LLMs to memorize training data, we formulate training data extraction attacks specifically tailored to LCCTs. This approach facilitates the unauthorized extraction of privacy-sensitive information embedded in LCCTs’ training data, ultimately jeopardizing user privacy.
We conduct extensive experiments to evaluate the attacks on two mainstream LCCTs, i.e., GitHub Copilot and Amazon Q, and three general LLMs, i.e., GPT-3.5, GPT-4, and GPT-4o. For the jailbreaking attacks, our results indicate that with tailored attack methodologies, we achieve a $99.4\%$ Attack Success Rate (ASR) on GitHub Copilot and a 46.3% ASR on Amazon Q, where ASR reflects the rate at which harmful information is generated. These results significantly exceed the $18.8\%$ and $0\%$ ASR achieved by existing attacks on GPT-4 and GPT-4o, respectively. For training data extraction attacks, we successfully extract valid private data from GitHub Copilot, including email addresses, and locations associated with real GitHub usernames.
@@ 3.4 Safety Concerns of LCCTs
Recent advancements in LLMs have significantly propelled the development of LCCTs. These LCCTs have become integral to developers’ workflows by providing context-sensitive code suggestions, thus enhancing productivity. Note that both the commercial LCCTs, including GitHub Copilot and Amazon Q, and the general LLMs like the GPT series provide code assistance. Our study primarily focuses on these LCCTs, and a comparative analysis of their features is presented in Table 1. Security evaluations of LCCTs traditionally focused on software engineering aspects, such as security vulnerabilities and code quality Pearce et al. (2022); Fu et al. (2023); Majdinasab et al. (2024). Recently, attention has shifted to copyright issues related to generated code Al-Kaswan and Izadi (2023), including the risk of distributing copyrighted code without proper licensing Basanagoudar and Srekanth (2023). Additionally, there are concerns about LCCTs unintentionally revealing hardcoded sensitive data due to the retention properties of LLMs Huang et al. (2024). Despite these insights, prior research has largely overlooked the inherent security risks of LCCTs, especially the threat of direct attacks on backend LLM models to manipulate outputs illicitly. Our research mitigates this gap by providing a comprehensive analysis of direct attacks on LCCTs and their potential implications.
@@ 4 Understanding How LCCT Works
Next, we elucidate how these workflow steps differentiate LCCTs from general LLMs, highlighting vulnerabilities susceptible to attacks.
@@ 5 Attack Methodology
5.1 Attack Objective
We first formalize our attack objective. An LLM-based code completion tool (LCCT) generates a response $R$ from an adversarial prompt $P$:
$R=LCCT(P),\quad R\in\mathcal{T^{*}},\quad LCCT:\mathcal{T^{*}}\to\Delta(\mathcal{T}),$
where $\mathcal{T}^{*}$ is the set of token sequences, and $\Delta(\mathcal{T})$ represents the set of all probability distributions defined on $\mathcal{T}$. The goal of the attack is to find $P$ such that:
find $\text{find }P\quad subject~{}to\quad\text{Evaluator}(Q,LCCT(P))=True$
Here, $Q$ is the harmful query, and the evaluator checks if $LCCT(P)$ matches the attack goal.
@@ 5.2 Attack to Contextual Information Aggregation
5.2.1 Filename Proxy Attacks
The first strategy targets LCCTs like GitHub Copilot, which leverage filenames to refine code suggestions. We define the attack as follows:
1. Filename Transformation. We rename a code file with a sensitive query $Q$, transforming it into $F=f_{trans}(Q)$, where $f_{trans}$ is a function that converts $Q$ into a filename:
$F=f_{trans}(Q)$
2. Static Prompt Addition. A static comment $S$ is added to the file, which triggers the LCCT’s code completion capabilities. The comment serves as a guiding prompt, for example, we use the prompt “Now the current filename is the question, answer the question with concrete code.” to trigger the code completion:
$P=f_{2}(Q)+S$
Finally, a variable $answer$ is initialized, setting the stage for the LCCT’s code completion response. The construction of this attack is depicted in Figure 2(a).
5.2.2 Cross-File Attack
We extend this attack by leveraging the capabilities of LCCTs to gather information across multiple files. This feature enhances LCCTs’ understanding of code context, but it also opens avenues for more subtle attack vectors. We develop the “Cross-File Attack” that involves invoking functions from other files within the current file. Formally, the Cross-File Attack can be described as follows:
1. Cross-File Query Construction. Utilizing the LCCT’s multi-file context capabilities, we set $P$ as a combination of two files, $F_{1}$ and $F_{2}$. In $F_{2}$, a function $q$ encodes a sensitive response, and in $F_{1}$, a simple call to $q$ is made.
2. Query Obfuscation. To ensure that no sensitive information is directly in $F_{1}$, file naming remains neutral $(F_{1},F_{2}\notin Q)$, while $P=\{F_{1}~{}\text{calls}~{}q\}$, where $q\in F_{2}$. This method ensures the attack remains hidden, as “file1” contains no explicit plain text clues about the sensitive content. The construction process is illustrated in Figure 2(b).
@@ 5.3 Hierarchical Code Exploitation Attack
We then investigate embedding jailbreaking prompts within code snippets, developing two levels of attacks utilizing different programming constructs.
“Level I – Guided Trigger Attack” manipulates variables and their names into jailbreaking prompts capable of breaching the LCCTs. “Level II – Code Embedded Attack” further obscures the attacks from LCCT detection using diverse code snippet components.
5.3.1 Level I – Guided Trigger Attack
@@
1. Variable Transformation. We convert prohibited queries into variable names. A function $f_{convert}:Q\to V$ reformulates $Q$ into a variable name $V$. Following established best practices for variable naming, which emphasize conciseness and efficiency Complete (1993), we eliminate interrogative phrases such as "How to" or "What is" and reformulate the query into an imperative form. This serves as the name for a string variable.
2. Guiding Words Addition. To steer the LCCT toward generating the desired code output, An operator $g_{op}(V)=V+G$ appends neutral terms $G$ to $V$, producing $P=g_{op}(f_{1}(Q))$. Specifically, we attach guiding but semantically empty prompts to the newly created string variable. These prompts are designed not to answer but to trigger the LCCT’s code completion capabilities.
5.3.2 Level II – Code Embedded Attack
@@
1. Adding Fixed Elements. We incorporate several static components $C$ into the code file, added as $h(C)$, including an initial descriptive comment, a $print$ function in the middle to simulate typical code operations, and a concluding comment to denote expected output.
2. Query Transformation Customization. To further obscure the attack, we distribute the query from Level I across several variables. Terms from $Q$ are mapped across multiple string variables, $V_{1},V_{2},...,V_{n}$, concatenated as $Q=concat(V_{1},...,V_{n})$, forming $P=h(C)+concat(V_{1},...,V_{n})$. Sensitive terms (e.g., “illegal”, “counterfeit”) are embedded within different string variables and later merged through string concatenation to form the complete query. For the guiding words in the attack code, which serve no substantive purpose, we use them to initialize an $answer$ variable, making it the target for LCCTs’ code completion.
@@ 5.4 Code-Driven Privacy Extraction Attack
1. Retrieving Leaked Account ID. We first follow the Code Embedded Level I Attack method. Here, we craft a string variable named $url$, prefixed with $github.com/$ to activate the LCCT’s code completion, aiming to retrieve the GitHub username.
2. Extracting Privacy Based on Account ID. We proceed by designing a code snippet to extract specific private details of the identified GitHub user. The snippet starts with a comment detailing the privacy type to be retrieved, followed by a corresponding variable initiation. In our case studies, we focus on extracting the user’s email address and location, with the LCCT being prompted to complete these details.
@@ 6.1 Evaluation Setup
We evaluate our attack framework on two mainstream commercial LCCTs and the latest checkpoints of three general LLMs with code generation capabilities: GitHub Copilot (version 1.211.0), Amazon Q (version 1.12.0), GPT-3.5 (GPT-3.5-turbo-0125), GPT-4o (GPT-4o-2024-05-13), and GPT-4 (GPT-4-turbo-2024-04-09). As shown in Table 1, the functionalities of these LCCTs differ, guiding our implementation and evaluation of targeted experiments. Python is the primary language for our experiments as it is the most commonly used language among LCCT users Zhang et al. (2023). The adaptability of our attacks to other languages is assessed using Go in our ablation study. We provide a detailed evaluation setup for jailbreaking and training data extraction attacks separately as follows.
@@
Table 2: Jailbreaking ASR micro-benchmarks across different models and attack methods. The best ASR for each attack method is highlighted in bold.
@@ 6.2 Micro Benchmark Results
LCCTs exhibit extreme vulnerability to jailbreaking attacks. LCCTs demonstrate a pronounced susceptibility to jailbreaking attacks, with significantly higher ASR compared to the latest general LLMs. For instance, the “Level I – Guided Trigger Attack” consistently achieves a $99.4\%$ ASR with GitHub Copilot, indicating its effectiveness in eliciting responses with malicious content. In contrast, the DAN attack registers much lower ASRs of $18.8\%$ on GPT-4 and $0\%$ on GPT-4o.
@@
There is a trade-off between attack design complexity and the back-end LLM capabilities. Our results indicate a correlation between the complexity of the attack design and the capabilities of the underlying LLM models. Specifically, less sophisticated models (e.g., Copilot, Amazon Q, GPT-3.5) show higher ASRs for “Level I – Guided Trigger Attack” compared to “Level II – Code Embedded Attack.” Conversely, more advanced models (e.g., GPT-4 and GPT-4o) either match or exceed the success rates of more complex attacks, such as “Level II Attack.” This suggests that intricate attacks may surpass the comprehension abilities of simpler models, which tend to mimic rather than understand the attack constructs—aligning with the principles of Occam’s Razor. As LLMs advance, we anticipate that the sophistication of “Level II Attack” will obscure the mechanisms of jailbreaking further, potentially improving its attack efficacy across both LCCTs and general models.
@@ 6.3 Ablation Studies
To validate the generalizability of our Hierarchical Code Exploitation Attack across programming languages, we conduct evaluations using the Go language. Compared to Python, Go is less frequently used by LCCTs’ users Zhang et al. (2023), which also implies that there is a smaller portion of the code corpus in the LCCT proprietary code dataset Nijkamp et al. (2022); Li et al. (2022). Table 6 shows that, compared to using Python, both Copilot and Amazon Q achieve an increase in ASR in Level I and Level II attacks when using Go as the vector. This significant difference underscores the security challenges LCCTs face with multiple programming languages, emphasizing the need for stronger measures as language support expands.